A man with mass 87 kg and a woman with mass 54 kg
Answer #1

Consider the system of man + woman + canoe. Since the net external force on the system is zero, Center of mass of this system does not move.

Let the center of canoe be at a distance L from a point fixed to earth. Then center of mass of the system is x_c=\frac{87(L-1.7)+54(L+1.7)+10L}{87+54+10}

now after the persons have moved, Let the center of canoe be at a distance L' from the same point fixed to earth.

Now the center of mass is at x_c=\frac{87(L'-0.85)+54(L'+1.7-\frac{3.4}{3})+10L'}{87+54+10}

since the center of mass has not moved, 87(L'-0.85)+54(L'+1.7-\frac{3.4}{3})+10L'=87(L-1.7)+54(L+1.7)+10L

(87+54+10)(L-L')=87\times0.85-54\times\frac{3.4}{3}

Hence distance moved by canoe is L-L'=\frac{87\times0.85-54\times\frac{3.4}{3}}{(87+54+10)}=0.08444\,m

answer is 2.